Earth and Mars are 1.0 AU and 1.52 AU away from the Sun, respectively. Given that the period of Earth's orbit around the Sun is 1 yr, calculate the period of the orbit of Mars around the Sun. (1 AU = 150 × 106 km)

a. 0.28 yr

b. 1.9 yr

c. 0.53 yr
Answer #1

R1= 1 AU and R2 =1.52 AU

T1 = 1 yr and T2 = ?

We know that , (T1^2)/(R1^3) = (T2^2)/(R2^3)

(1^2)/(1^3) =(T2^2)/(1.52^3)

1*3.512 = T2^2

T2 = 1.874 AU

Approximately , T2 = 1.9 AU